# von Neumann Orbit Equivalence

### v1

@@ 2.4. Modules over tracial von Neumann algebras
    cyclic $\overline{\mathrm{Span}\{x\xi y:x\in A,y\in B\}}=\mathcal{H}$.
@@ 3. Von Neumann Orbit Equivalence
In this section, we define von Neumann orbit equivalence for tracial von Neumann algebras and for countable discrete groups. We shall see that groups are von Neumann orbit equivalent if and only if the corresponding group von Neumann algebras are, and we conclude this section with the proof that von Neumann orbit equivalent tracial von Neumann algebras are von Neumann equivalent in the sense of [IPR24]
@@ 3.1. Von Neumann orbit equivalence for tracial von Neumann algebras
Let $(A,\tau_{A})$, and $(B,\tau_{B})$ be tracial von Neumann algebras. We say that $A$ is von Neumann orbit equivalent to $B$, denoted $A\sim_{\mathrm{vNOE}}B$, if either of the two equivalent conditions in Theorem 3.1 is satisfied. If $A$ is von Neumann orbit equivalent to $B$, then the triple $(\mathcal{H},Q,\xi)$ or the pair $(Q,\phi)$ of Theorem 3.1 will be called a vNOE-coupling between $A$ and $B$.
@@
Let $A_{i},B_{i},i=1,2$ be tracial von Neumann algebras such that $A_{i}\sim_{\mathrm{vNOE}}B_{i},i=1,2$. Then, $A_{1}*A_{2}\sim_{\mathrm{vNOE}}B_{1}*B_{2}$.
@@ 3.3. Relationship to von Neumann equivalence
4. Graph product and von Neumann orbit equivalence
We now turn to prove Theorem 1.7, and prove the following stronger Theorem.
@@ 4. Graph product and von Neumann orbit equivalence
Let $\mathcal{G}=(V,E)$ be a simple finite graph. Let $A$ and $B$ be the graph products over $\mathcal{G}$ of tracial vertex von Neumann algebras $\{(A_{v},\tau_{A_{v}}):v\in V\}$ and $\{(B_{v},\tau_{B_{v}}):v\in V\}$, respectively. For a subset $U\subseteq V$, we set $A_{U}=\bigstar_{v\in U}A_{v}$, and similarly, $B_{U}=\bigstar_{v\in U}B_{v}$. There exist a tracial von Neumann algebra $(Q,\tau_{Q})$ and a *-homomorphism $\phi:A\to B\otimes Q$ such that the following hold:
  1. (1)
    $\phi$ satisfies condition (2) of Theorem 3.1;
  2. (2)
    $\phi(A_{U})\subseteq B_{U}\,\overline{\otimes}\,Q$ for any subset $U\subseteq V$; and
  3. (3)
    $\overline{\operatorname{Span}\{\phi(y)x:y\in A_{U},x\in Q\}}^{\|\cdot\|_{2}}=L^{2}(B_{U}\,\overline{\otimes}\,Q)$ for any subset $U\subseteq V$.
In particular, $A$ is von Neumann orbit equivalent to $B$.
Remark 4.2.
Note that condition (2) of the theorem automatically implies that ${\mathbb{E}}_{Q,U}\circ\phi=\tau_{A_{U}}$, where ${\mathbb{E}}_{Q,U}$ is the conditional expectation from $B_{U}\,\overline{\otimes}\,Q$ onto $Q$. Indeed, if ${\mathbb{E}}_{Q}$ is the conditional expectation from $B\,\overline{\otimes}\,Q$ onto $Q$, then ${\mathbb{E}}_{Q,U}={\mathbb{E}}_{Q}|_{B_{U}\,\overline{\otimes}\,Q}$, $\tau_{A_{U}}=\tau_{A}|_{A_{U}}$, and ${\mathbb{E}}_{Q}\circ\phi=\tau_{A}$. In particular, (2) and (3) together imply that the pair $(Q,\phi)$ is a vNOE coupling between $A_{U}$ and $B_{U}$ for any subset $U\subseteq V$.
Proof of Theorem 4.1.
We proceed by induction on the number of vertices in the graph. The statement is true for a graph with only one vertex. Suppose the theorem is true for graphs with $|V|-1$ vertices. Fix a vertex $v\in V$ and set $V^{\prime}=V\setminus\{v\}$. Since $A_{v}\sim_{\mathrm{vNOE}}B_{v}$, there exist a tracial von Neumann algebra $(Q_{1},\tau_{Q_{1}})$ and a *-homomorphism $\phi_{1}:A_{v}\to B_{v}\,\overline{\otimes}\,Q_{1}$ satisfying condition (2) in Theorem 3.1. Moreover, by the induction hypothesis, we also have a tracial von Neumann algebra $(Q_{2},\tau_{Q_{2}})$ and a *-homomorphism $\phi_{2}:A_{V^{\prime}}\to B_{V^{\prime}}\,\overline{\otimes}\,Q_{2}$ satisfying (1)-(3) in the statement of the theorem. In particular, we have that $\phi_{2}(A_{\mathsf{lk}(v)})\subseteq B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q_{2}$ and that $\overline{\operatorname{Span}\{\phi_{2}(y)x:y\in A_{\mathsf{lk}(v)},x\in Q_{2}\}}^{\|\cdot\|_{2}}=L^{2}(B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q_{2}).$
Recall that we have the following decompositions of $A$ and $B$ as amalgamated free products:
$A=(A_{v}\,\overline{\otimes}\,A_{\mathsf{lk}(v)})*_{A_{\mathsf{lk}(v)}}A_{V^{\prime}},\qquad B=(B_{v}\,\overline{\otimes}\,B_{\mathsf{lk}(v)})*_{B_{\mathsf{lk}(v)}}B_{V^{\prime}},\qquad$
Set $Q=Q_{1}\,\overline{\otimes}\,Q_{2}$ and let $\tau_{Q}$ be the trace on $Q$. We now construct a *-homomorphism
$\phi:(A_{v}\,\overline{\otimes}\,A_{\mathsf{lk}(v)})*_{A_{\mathsf{lk}(v)}}A_{V^{\prime}}\to(B_{v}\,\overline{\otimes}\,B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q)*_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}(B_{V^{\prime}}\,\overline{\otimes}\,Q).$
For $x\in A_{v},y\in A_{\mathsf{lk}(v)}$, and $z\in A_{V^{\prime}}$, we define
$\phi(x\otimes y)$ $=(\phi_{1}(x)\otimes\phi_{2}(y))*1,$
$\phi(z)$ $=1*\tilde{\phi}_{2}(z),$
where $\tilde{\phi}_{2}:A_{V^{\prime}}\to B_{V^{\prime}}\,\overline{\otimes}\,Q_{2}\,\overline{\otimes}\,Q_{1}$ is given by $\tilde{\phi}_{2}(z)=\phi_{2}(z)\otimes 1_{Q_{1}},z\in A_{V^{\prime}}$. Note that $\phi$ is well-defined since $\phi_{2}(A_{U^{\prime}})\subseteq B_{U^{\prime}}\,\overline{\otimes}\,Q_{2}$, and hence $\tilde{\phi}_{2}(A_{U^{\prime}})\subseteq B_{U^{\prime}}\,\overline{\otimes}\,Q$ for any subset $U^{\prime}\subseteq V^{\prime}$.
Let ${\mathbb{E}}_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}:B\,\overline{\otimes}\,Q\to B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q$ and ${\mathbb{E}}_{A_{\mathsf{lk}(v)}}:A\to A_{\mathsf{lk}(v)}$ be the normal conditional expectations. We show that ${\mathbb{E}}_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}\circ\phi=\phi\circ{\mathbb{E}}_{A_{\mathsf{lk}(v)}}$ by showing that an alternating centered word with respect to ${\mathbb{E}}_{A_{\mathsf{lk}(v)}}$ gets mapped to an alternating centered word with respect to ${\mathbb{E}}_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}$ under $\phi$. Let $x\otimes y\in A_{v}\,\overline{\otimes}\,A_{\mathsf{lk}(v)}$ be a simple tensor such that ${\mathbb{E}}_{A_{\mathsf{lk}(v)}}(x\otimes y)=0$. Since ${\mathbb{E}}_{A_{\mathsf{lk}(v)}}|_{A_{v}\,\overline{\otimes}\,A_{\mathsf{lk}(v)}}=\tau_{A_{v}}\otimes\operatorname{id}_{A_{\mathsf{lk}(v)}}$, it follows that either $\tau_{A_{v}}(x)=0$ or $y=0$. If ${\mathbb{E}}_{Q_{1}}:B_{v}\,\overline{\otimes}\,Q_{1}\to Q_{1}$ is the normal conditional expectation, then we note that ${\mathbb{E}}_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}|_{B_{v}\,\overline{\otimes}\,B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}={\mathbb{E}}_{Q_{1}}\otimes\operatorname{id}_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q_{2}}$. Since ${\mathbb{E}}_{Q_{1}}\circ\phi_{1}=\tau_{A_{v}}$, we have that
$({\mathbb{E}}_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}\circ\phi)(x\otimes y)$ $={\mathbb{E}}_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}(\phi_{1}(x)\otimes\phi_{2}(y))$
$=({\mathbb{E}}_{Q_{1}}\otimes\operatorname{id}_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q_{2}})(\phi_{1}(x)\otimes\phi_{2}(y))$
$={\mathbb{E}}_{Q_{1}}(\phi_{1}(x))\otimes\phi_{2}(y)$
$=\tau_{A_{v}}(x)\otimes\phi_{2}(y)=0.$
Next, suppose we have $z\in A_{V^{\prime}}$ such that ${\mathbb{E}}_{A_{\mathsf{lk}(v)}}(z)=0$. Since ${\mathbb{E}}_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}$ is trace-preserving and $B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q$-bimodular, so, to show that ${\mathbb{E}}_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}(\phi(z))=0$, it suffices to show that $\tau_{B_{V^{\prime}}\,\overline{\otimes}\,Q}(w\tilde{\phi}_{2}(z))=0$ for all $w\in B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q$. Furthermore, by induction hypothesis, we have that $L^{2}(B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q_{2})=\overline{\operatorname{Span}\{\phi_{2}(x)y:x\in A_{\mathsf{lk}(v)},y\in Q_{2}\}}$. Therefore, we get that
$L^{2}(B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q)=\overline{\operatorname{Span}\{\tilde{\phi}_{2}(x)(y_{1}\otimes y_{2}):x\in A_{\mathsf{lk}(v)},y_{1}\in Q_{1},y_{2}\in Q_{2}\}}.$
Thus, it suffices to show that $\tau_{B_{V^{\prime}}\,\overline{\otimes}\,Q}(\tilde{\phi}_{2}(x)\tilde{\phi}_{2}(z)(y_{1}\otimes y_{2}))=0$ for every $y_{1}\in Q_{1},y_{2}\in Q_{2},x\in A_{\mathsf{lk}(v)}$. If ${\mathbb{E}}_{Q_{2}}:B_{V^{\prime}}\,\overline{\otimes}\,Q_{2}\to Q_{2}$ and ${\mathbb{E}}_{Q}:B_{V^{\prime}}\,\overline{\otimes}\,Q\to Q$ are normal conditional expectations, then we have that ${\mathbb{E}}_{Q}=\operatorname{id}_{Q_{1}}\otimes{\mathbb{E}}_{Q_{2}}$, and moreover, since $\tau_{Q}\circ{\mathbb{E}}_{Q}=\tau_{B_{V^{\prime}}\,\overline{\otimes}\,Q}$, and $\tau_{Q_{2}}\circ{\mathbb{E}}_{Q_{2}}=\tau_{A_{V^{\prime}}}$ (by induction hypothesis), we get that
$\tau_{B_{V^{\prime}}\,\overline{\otimes}\,Q}(\tilde{\phi}_{2}(xz)(y_{1}\otimes y_{2}))$ $=\tau_{Q}((\operatorname{id}_{Q_{1}}\otimes{\mathbb{E}}_{Q_{2}})(1_{Q_{1}}\otimes\phi_{2}(xz))(y_{1}\otimes y_{2}))$
$=\tau_{Q_{1}}(y_{1})\tau_{Q_{2}}(y_{2}{\mathbb{E}}_{Q_{2}}(\phi_{2}(xz))$
$=\tau_{Q_{1}}(y_{1})\tau_{Q_{2}}(\tau_{A_{V^{\prime}}}(xz)y_{2})$
$=\tau_{Q}(y_{1}\otimes y_{2})\tau_{A_{\mathsf{lk}(v)}}({\mathbb{E}}_{A_{\mathsf{lk}(v)}}|_{A_{V^{\prime}}}(xz))$
$=\tau_{Q}(y_{1}\otimes y_{2})\tau_{A_{\mathsf{lk}(v)}}(x{\mathbb{E}}_{A_{\mathsf{lk}(v)}}(z))=0.$
Therefore, $\phi$ maps an alternating centered word with respect to ${\mathbb{E}}_{A_{\mathsf{lk}}(v)}$ to an alternating centered word with respect to ${\mathbb{E}}_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}$, and thus we have that ${\mathbb{E}}_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}\circ\phi=\phi\circ{\mathbb{E}}_{A_{\mathsf{lk}(v)}}.$ Therefore, if $\widetilde{{\mathbb{E}}}_{Q}$ is the normal conditional expectation from $B\,\overline{\otimes}\,Q\to Q$, then we have
$\widetilde{{\mathbb{E}}}_{Q}\circ\phi$ $={\mathbb{E}}_{Q}|_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}\circ{\mathbb{E}}_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}\circ\phi$
$={\mathbb{E}}_{Q}|_{B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q}\circ\phi\circ{\mathbb{E}}_{A_{\mathsf{lk}(v)}}$
$=\tau_{A_{\mathsf{lk}(v)}}\circ{\mathbb{E}}_{A_{\mathsf{lk}(v)}}$
$=\tau_{A}.$
Next, we show that
$\overline{\operatorname{Span}\{\phi(x)y:x\in(A_{v}\,\overline{\otimes}\,A_{\mathsf{lk}(v)})*_{A_{\mathsf{lk}(v)}}A_{V^{\prime}},y\in Q\}}^{\|\cdot\|_{2}}=L^{2}(((B_{v}\,\overline{\otimes}\,B_{\mathsf{lk}(v)})*_{B_{\mathsf{lk}(v)}}B_{V^{\prime}})\,\overline{\otimes}\,Q).$
To this end, set $W=\overline{\operatorname{Span}\{\phi(x)y:x\in(A_{v}\,\overline{\otimes}\,A_{\mathsf{lk}(v)})*_{A_{\mathsf{lk}(v)}}A_{V^{\prime}},y\in Q\}}^{\|\cdot\|_{2}}$. Since $W$ is invariant under right multiplication by elements in $Q$, it suffices to show that $((B_{v}\,\overline{\otimes}\,B_{\mathsf{lk}(v)})*_{B_{\mathsf{lk}(v)}}B_{V^{\prime}})\otimes 1_{Q_{1}}\subseteq W$. Furthermore, by Lemma 3.5, it suffices to show that $W$ contains $B_{v}\otimes 1,1\otimes B_{\mathsf{lk}(v)}$, and $B_{V^{\prime}}\otimes 1$. Since $\phi|_{A_{v}}=\phi_{1},\phi|_{A_{\mathsf{lk}(v)}}=\phi_{2}$, and $\phi|_{A_{V^{\prime}}}=\tilde{\phi}_{2}$, we have the following:
@@
$W$ $\supseteq\overline{\operatorname{Span}\{\phi_{1}(x)y:x\in A_{v},y\in Q_{1}\}}^{\|\cdot\|_{2}}=L^{2}(B_{v}\,\overline{\otimes}\,Q_{1})$
$W$ $\supseteq\overline{\operatorname{Span}\{\phi_{2}(x)y:x\in A_{\mathsf{lk}(v)},y\in Q_{2}\}}^{\|\cdot\|_{2}}=L^{2}(B_{\mathsf{lk}(v)}\,\overline{\otimes}\,Q_{2})$
$W$ $\supseteq\overline{\operatorname{Span}\{\tilde{\phi}_{2}(x)y:x\in A_{V^{\prime}},y\in Q_{2}\}}^{\|\cdot\|_{2}}=L^{2}(B_{V^{\prime}}\,\overline{\otimes}\,Q_{2}).$
Thus, condition (1) in the statement of the theorem holds. Finally, we show that conditions (2) and (3) hold. For this, let $U\subseteq V$ be a subset of vertices and first suppose that $v\notin U$. Then, $U\subseteq V^{\prime}$ and $A_{U}\subseteq A_{V^{\prime}}$, whence it follows that $\phi|_{A_{U}}=\tilde{\phi}_{2}|_{A_{U}}$. Thus, (2) and (3) hold for $U$. Now suppose $v\in U\subseteq V$. Then, set $U^{\prime}=U\setminus\{v\}\subseteq V^{\prime}$, and we have that
$A_{U}=(A_{v}\,\overline{\otimes}\,A_{\mathsf{lk}(v)\cap U})*_{A_{\mathsf{lk}(v)\cap U}}A_{U^{\prime}},\quad B_{U}\,\overline{\otimes}\,Q=(B_{v}\,\overline{\otimes}\,B_{\mathsf{lk}(v)\cap U}\,\overline{\otimes}\,Q)*_{B_{\mathsf{lk}(v)\cap U}\,\overline{\otimes}\,Q}(B_{U^{\prime}}\,\overline{\otimes}\,Q).$
Now, since $U^{\prime}\subset V^{\prime}$, so, by the induction hypothesis, the pair $(Q_{2},\phi_{2})$ satisfies conditions (2) and (3) for any subset $U^{\prime\prime}\subseteq U^{\prime}$. Using the vNOE coupling $(Q_{1},\phi_{1})$ between $A_{v}$ and $B_{v}$ and the pair $(\phi_{2},Q_{2})$, if we repeat the same construction as above to construct a *-homomorphism $\psi:A_{U}\to B_{U}\otimes Q$, we obtain that $\psi=\phi|_{A_{U}}$ and that the pair $(Q,\psi)$ is a vNOE coupling between $A_{U}$ and $B_{U}$. Thus, conditions (2) and (3) hold for $U$ and this finishes the proof.

### v2

@@ 2.2. Free product and amalgamated free product
2.3. Graph product
Let $\mathcal{G}=(V,E)$ be a simple graph with the vertex set $V$ and the edge set $E\subseteq V\times V\setminus\{(v,v):v\in V\}$. We assume that the graph $\mathcal{G}$ is non-oriented, i.e., $(v,w)\in E$ if and only if $(w,v)\in E$. A word $v_{1}v_{2}\cdots v_{n}$ of vertices in $V$ is called reduced if it satisfies the following property: if there exist $k<l$ such that $v_{k}=v_{l}$, then there is some $k<j<l$ such that $(v_{k},v_{j})\notin E$. Let $\mathcal{G}=(V,E)$ be a simple graph, $(A,\tau)$ be a tracial von Neumann algebra, and $\{(A_{v},\tau_{v}):v\in V\}$ be a family of tracial von Neumann subalgebras of $(A,\tau)$ such that $\tau|_{A_{v}}=\tau_{v}$ for all $v\in V$. We say that the family $\{(A_{v},\tau_{v}):v\in V\}$ is $\mathcal{G}$-independent if the following property holds: if $v_{1}\cdots v_{n}$ is a reduced word and $a_{1},\ldots,a_{n}\in A$ are such that $a_{i}\in A_{v_{i}}$ and $\tau(a_{i})=0$, then $\tau(a_{1}\cdots a_{n})=0$. On the other hand, given a simple graph $\mathcal{G}=(V,E)$ and a family of tracial von Neumann algebras $\{(A_{v},\tau_{v}):v\in V\}$, there is a unique, up to isomorphism, tracial von Neumann algebra $(A,\tau)$, called the graph product over $\mathcal{G}$ of the family $\{(A_{v},\tau_{v}):v\in V\}$, and trace-preserving inclusions $\varphi_{v}:A_{v}\hookrightarrow A$ such that the family $\{\varphi_{v}(A_{v}):v\in V\}$ is $\mathcal{G}$-independent and generates $A$ as a von Neumann algebra (see [Mło04, CF17]). We denote the graph product $(A,\tau)$ of the family $\{(A_{v},\tau_{v}):v\in V\}$ by
$(A,\tau)=\bigstar_{v\in V}(A_{v},\tau_{v}).$
Remark 2.1.
@@ 2.4. Modules over tracial von Neumann algebras
    cyclic if $\overline{\mathrm{Span}\{x\xi y:x\in A,y\in B\}}=\mathcal{H}$.
@@ 3. Von Neumann Orbit Equivalence
In this section, we define von Neumann orbit equivalence for tracial von Neumann algebras and for countable discrete groups. We shall see that groups are von Neumann orbit equivalent if and only if the corresponding group von Neumann algebras are, and we conclude this section with the proof that von Neumann orbit equivalent tracial von Neumann algebras are von Neumann equivalent in the sense of [IPR24].
@@ 3.1. Von Neumann orbit equivalence for tracial von Neumann algebras
Let $(A,\tau_{A})$, and $(B,\tau_{B})$ be tracial von Neumann algebras. We say that $(A,\tau_{A})$ is von Neumann orbit equivalent to $(B,\tau_{B})$, denoted $(A,\tau_{A})\sim_{\mathrm{vNOE}}(B,\tau_{B})$, if either of the two equivalent conditions in Theorem 3.1 is satisfied. If $(A,\tau_{A})$ is von Neumann orbit equivalent to $(B,\tau_{B})$, then the triple $(\mathcal{H},Q,\xi)$ or the pair $(Q,\phi)$ of Theorem 3.1 will be called a vNOE-coupling between $(A,\tau_{A})$ and $(B,\tau_{B})$.
Remark 3.4.
As mentioned in the introduction, the above definition depends on the choice of the traces $\tau_{A}$ and $\tau_{B}$. When in a situation where the traces are fixed and there is no risk of confusion, we will often drop them, and simply say that $A$ and $B$ are vNOE and write $A\sim_{\mathrm{vNOE}}B$.
@@
If $(A_{i},\tau_{A_{i}}),(B_{i},\tau_{B_{i}}),~{}i=1,2$ are tracial von Neumann algebras such that $(A_{i},\tau_{A_{i}})\sim_{\mathrm{vNOE}}(B_{i},\tau_{B_{i}}),~{}i=1,2$, then, $(A_{1}*A_{2},\tau_{A_{1}}*\tau_{A_{2}})\sim_{\mathrm{vNOE}}(B_{1}*B_{2},\tau_{B_{1}}*\tau_{B_{2}})$.
@@ 3.3. Relationship to von Neumann equivalence
We conclude this section by proving Theorem 1.8, which we recall below.
@@
Let $\mathcal{G}=(V,E)$ be a simple graph, with at most countably infinite vertices. Let $(A,\tau_{A})$ and $(B,\tau_{B})$ be two graph products over $\mathcal{G}$, with tracial vertex von Neumann algebras $\{(A_{v},\tau_{A_{v}})\}_{v\in V}$ and $\{(B_{v},\tau_{B_{v}})\}_{v\in V}$, respectively. If $(A_{v},\tau_{A_{v}})\sim_{\mathrm{vNOE}}(B_{v},\tau_{B_{v}})$ for every $v\in V$, then $(A,\tau_{A})\sim_{\mathrm{vNOE}}(B,\tau_{B})$.
Proof.
Let $(Q_{v},\phi_{v})$ be a vNOE-coupling between $(A,\tau_{A_{v}})$ and $(B_{v},\tau_{B_{v}})$, where $\phi_{v}:A_{v}\to B_{v}\,\overline{\otimes}\,Q_{v}$ is a normal unital *-homomorphism satisfying
@@
${\mathbb{E}}_{v}\circ\phi_{v}=\tau_{A_{v}}\qquad\text{ and }\qquad\overline{\operatorname{Span}\{\phi_{v}(a)x:a\in A_{v},x\in Q_{v}\}}^{\|\cdot\|_{2}}=L^{2}(B_{v}\,\overline{\otimes}\,Q_{v}),$
here, ${\mathbb{E}}_{v}:B_{v}\,\overline{\otimes}\,Q_{v}\to Q_{v}$ is the normal conditional expectation. Define $(Q,\tau_{Q})$ to be the tensor product ¯$\,\overline{\otimes}\,_{v\in V}(Q_{v},\tau_{Q_{v}})$. We now define a *-homomorphism $\phi:A\to B\,\overline{\otimes}\,Q$ and show that it has the desired properties. To this end, first note that if $v$ and $w$ are two distinct vertices connected by an edge, then $B_{v}$ commutes with $B_{w}$ and $Q_{v}$ commutes with $Q_{w}$ as $v\neq w$. Thus, $\phi_{v}(A_{v})$ and $\phi_{w}(A_{w})$ commute inside $B\,\overline{\otimes}\,Q$, and therefore, we can define $\phi$ on the algebraic graph product (i.e., the universal unital *-algebra generated by the $\{A_{v}\}_{v\in V}$ subject to the relation that $A_{v}$ commutes with $A_{w}$ whenever $v$ and $w$ are connected by an edge). In particular, note that $\phi(a)=\phi_{v}(a)$, whenever $a\in A_{v}$. Next, we show that $\phi$ extends to a normal unital *-homomorphism on $A$ and satisfies ${\mathbb{E}}_{Q}(\phi(a))=\tau_{A}(a)$ for all $a\in A$, where ${\mathbb{E}}_{Q}:B\,\overline{\otimes}\,Q\to Q$ is the normal conditional expectation. To see this, we note that for any reduced word $v=v_{1}\cdots v_{n}$ of vertices, and elements $a_{i}\in A_{v_{i}}$ of trace zero, $1\leq i\leq n$, we have that
${\mathbb{E}}_{Q}(\phi_{1}(a_{1})\cdots\phi_{n}(a_{n}))=0.$
Indeed, since ${\mathbb{E}}_{v_{i}}(\phi_{v_{i}}(a_{v_{i}}))=\tau_{A_{v_{i}}}(a_{v_{i}})=0,~{}1\leq i\leq n$, so, $\phi_{v_{i}}(a_{v_{i}})$ can be approximated by linear combinations of elements of the form $b_{v_{i}}\otimes q_{v_{i}}$, where $b_{v_{i}}\in B_{v_{i}},q_{v_{i}}\in Q_{v_{i}}$, and $\tau_{B_{v_{i}}}(b_{v_{i}})=0,~{}1\leq i\leq n$. Thus, ${\mathbb{E}}_{Q}(\phi_{1}(a_{1})\cdots\phi_{n}(a_{n}))=0$ and hence it follows that the map $\phi$, defined on the algebraic graph product, is trace-preserving. Therefore, by [CF17, Proposition 3.22], $\phi$ extends to a well-defined normal unital *-homomorphism on $A$ and satisfies ${\mathbb{E}}_{Q}(\phi(a))=\tau_{A}(a)$ for all $a\in A$. Finally, to show that $\overline{\operatorname{Span}\{\phi(a)x:a\in A,x\in Q\}}^{\|\cdot\|_{2}}=L^{2}(B\,\overline{\otimes}\,Q)$, it suffices to show that
(*) $\overline{\operatorname{Span}(\phi_{v_{1}}(A_{v_{1}})\cdots\phi_{n}(A_{v_{n}})(1\otimes Q))}^{\|\cdot\|_{2}}=\overline{\operatorname{Span}(B_{v_{1}}\cdots B_{v_{n}}(1\otimes Q))}^{\|\cdot\|_{2}},$
for any reduced word $v=v_{1}\cdots v_{n}$ of vertices and for all $n$. We proceed by induction. It is straightforward to see that for any vertex $v$, $\overline{\operatorname{Span}(\phi_{v}(A_{v})(1\otimes Q))}^{\|\cdot\|_{2}}=L^{2}(B_{v}\,\overline{\otimes}\,Q)=\overline{\operatorname{Span}(B_{v}(1\otimes Q))}^{\|\cdot\|_{2}}$. Suppose that (*3.3) holds for any reduced word $v_{1}\cdots v_{n}$ of length $n$ and let $v_{1}\cdots v_{n}v_{n+1}$ be any reduced word of length $n+1$. Then,
$\overline{\operatorname{Span}(\phi_{v_{1}}(A_{v_{1}})\cdots\phi_{n+1}(A_{n+1})(1\otimes Q))}^{\|\cdot\|_{2}}$ $=\overline{\operatorname{Span}(\phi_{v_{1}}(A_{v_{1}})B_{v_{2}}\cdots B_{v_{n+1}}(1\otimes Q))}^{\|\cdot\|_{2}}$
$=\overline{\operatorname{Span}(\phi_{v_{1}}(A_{v_{1}})(1\otimes Q)(B_{v_{2}}\cdots B_{v_{n+1}}\otimes 1))}^{\|\cdot\|_{2}}$
$=\overline{\operatorname{Span}(B_{v_{1}}(1\otimes Q)(B_{v_{2}}\cdots B_{v_{n+1}}\otimes 1))}^{\|\cdot\|_{2}}$
$=\overline{\operatorname{Span}(B_{v_{1}}B_{v_{2}}\cdots B_{v_{n+1}}(1\otimes Q))}^{\|\cdot\|_{2}}.$
Thus, $(A,\tau_{A})\sim_{\mathrm{vNOE}}(B,\tau_{B})$. ∎